OpenStudy (daenio):

There are 15 players on a co-ed volleyball team. There are 10 girls and 5 boys. The coach would like to send 4 players to a weekend volleyball clinic. a) What is the probability that at least 1 of the players sent to the clinic is a girl? [3] P (x=1)= (9 C 1) (6 C 5) / 15 C 4 = 9 x 6/ 1365 = 0.040 The probability is 0.040. b) What is the expected number of girls to be sent to the clinic? [2] E(x)= 4 x 10/ 15 = 2.667

OpenStudy (daenio):

Hey, I was wondering if I did this right. Could you check? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

a) P(No girls sent) = ((10 C 0)*(5 C 4))/(15 C 4) P(No girls sent) = (1*5)/(1365) P(No girls sent) = (5)/(1365) P(No girls sent) = 1/273 P(At least one girl sent) = 1 - P(No girls sent) P(At least one girl sent) = 1 - 1/273 P(At least one girl sent) = 272/273 P(At least one girl sent) = 0.996336996

OpenStudy (daenio):

Thanks again!

OpenStudy (daenio):

I was way off then, lol.

jimthompson5910 (jim_thompson5910):

one sec while I work the other

jimthompson5910 (jim_thompson5910):

just wanted to confirm the answer with a simulation, but you have the correct answer for part b), nice job.

OpenStudy (daenio):

Cool, thanks man. :)

jimthompson5910 (jim_thompson5910):

yw